General Electric has just issued a callable​ (at par)​ 10-year, 6.1 % coupon bond with annual coupon payments. The bond can be called at par in one year or anytime thereafter on a coupon payment date. It has a price of $ 101.64. a. What is the​ bond's yield to​ maturity? b. What is its yield to​ call? c. What is its yield to​ worst?
Answer #1

1.
=RATE(10,6.1%*1000,-1016.4,1000)=5.87845787215695%

2.
=RATE(1,6.1%*1000,-1016.4,1000)=4.38803620621801%

3.
=RATE(1,6.1%*1000,-1016.4,1000)=4.38803620621801%